# Membership Inference Attacks on Large-Scale Models: A Survey

### v1

@@ Abstract
The adoption of the Large Language Model (LLM) has accelerated dramatically since the ChatGPT from OpenAI went online in November 2022. Recent advances in Large Multimodal Models (LMMs), which process diverse data types and enable interaction through various channels, have expanded beyond the text-to-text limitations of early LLMs, attracting significant and concurrent attention from both researchers and industry. While LLMs and LMMs are starting to spread widely, concerns about their privacy risks are increasing as well. Membership Inference Attacks (MIAs), techniques used to determine whether a particular data point was part of a model’s training set, serve as a key metric for assessing the privacy vulnerabilities of machine learning models. Hu et al. show that various machine learning algorithms are vulnerable to MIA. Despite extensive studies on MIAs in traditional models, there remains a lack of systematic surveys addressing their effectiveness and implications in modern large-scale models like LLMs and LMMs. In this paper, we systematically reviewed recent studies of MIA against LLMs and LMMs. We analyzed and categorized each attack based on their methodology and scenario and discussed the limitations in existing research. Additionally, we examine privacy concerns associated with the fine-tuning process. Finally, we provided some suggestions for future research in this direction.
@@ 1 Introduction
As LLMs and LMMs continue to evolve and gain widespread adoption, addressing their privacy concerns has become increasingly critical. In the LLM area, various research studies have already been conducted that focus on the general privacy risks of LLM [17, 18]. However, these studies provide limited depth and focus in their analysis of MIA. In the field of multimodal models, apart from the work of M.A. Rahman et al. [19], which examines privacy concerns from an educational perspective, and a general review by S.K. Tetarave et al. [20], there remains a lack of in-depth and systematic analysis. Regarding MIA, the study conducted by H. Hu et al. [21] has summarized MIA against various basic machine learning models. While their survey provides a valuable reference for researchers in the area, the study is up to 2022, and there is a lack of study about MIA on more advanced machine learning models. A similar study was published in 2024 by J. Niu et al. [22], which still only summarizes the research attack up to Jan. 2024 and lacks analysis and focus on advanced models. Table 1 summarizes the limitations of existing surveys and highlights the contributions of our research.
@@
2. An analysis of Fine-Tune Database Membership Inference Attack in advanced machine learning models. Given that fine-tuning technology is widely used in deploying LLMs and LMMs, we refine the attack setting for fine-tuning scenarios and review existing MIA research that targets the fine-tuned database’s exposure.
3. Suggestions for future research. Based on the challenges and research gap identified, we provide several suggestions for further exploration in MIA against modern AI models.
@@
Research Large-Scale Model MIA FT Attack Survey Up
LLM LMM In-Depth In-Depth To
@@ 2.1 Adversarial knowledge
Table 2: Summarize of attack scenario
@@
2.2 Attack strategies
3 Membership Inference Attack against Large Language Model
3.1 Black-Box & Target Model Based
MIN-KPROB [30] The MIN-KPROB MIA is based on the assumption that the response from non-member datapoint is more likely to have the words in the sentence that are less like the words in the rest of the sentence, which are called “outlier,” compared to the member example. An outlier is determined by tokenizing the whole sentence $x$ and calculating the log-likelihood of its specific token $x_{i}$ to the rest of the sentence. To predict the membership identity of a datapoint, it first selects kof tokens that have minimal log-likelihood from the sentence, denote as $x_{MIN-K\%}=(x_{1},x_{2},\cdots,x_{N})$, and calculates the average of the log-likelihood of this set. The k is a pre-designed number. The following equation shows the formula of the result of MIN-KPROB MIA:
$R=\frac{1}{N}\sum_{i=1}^{N}logp(x_{i}|x)$
The MIA model will claim membership identity if the result is higher than a pre-designed threshold. However, its accuracy depends on the pre-designed K and threshold value, which made this method less practical.
Perplexity-Based MIA [31, 32] This MIA method follows the idea that machine learning models tend to have better accuracy when predicting membership identity. It analyzes the average of the perplexity of the whole sentence $x=(x_{1},x_{2},…,x_{N})$, given by:
$exp\left(-\frac{1}{N}\sum_{i=1}^{N}logf_{\theta}(x_{i}|x_{1},x_{2},\cdots,x_{i})\right)$
The MIA model will accept that a data point is a member if the perplexity is lower than a pre-designed threshold, which means the target model is more confident with its prediction. It has several variances including loss based-MIA and zlib entropy based-MIA. However, this model is safer from a high false-positive rate since a well-optimized LLM is able to achieve a high accuracy even for unseen data.
Sensitivity-Based MIA [33, 34, 35] The sensitivity-based MIA model predicts the membership identity of the input data based on the robustness of the target model’s prediction. Specifically, given an input x to be inferred by the target model, the adversary first generates a reference dataset $x^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\cdots,x_{N}^{\prime})$ by adding perturbation, including randomly deleting words, changing the case of the character, or adding space inside sentences. The target model $f_{target}(x)$ will infer both the original data and the data from the reference dataset, and the prediction is based on the difference in the performance of the target model between the original and the modified data. A threshold is pre-defined and if the result passes the threshold, the attack model predicts it is from the training dataset.
$R=\frac{1}{N}\sum_{i=1}^{N}(f_{t}(x)-f_{t}(x_{i}^{\prime}))$
This attack method assumes the machine learning model is specifically more confident in the prediction of member data and, therefore, becomes more sensitive and yields a larger gap between the original and the reference data. The limitation of this method is that the threshold of the gap needs to be hand-designed, and it’s hard to maintain the consistency of the perturbation added, which could impact the detection result.
MIN-K%++ MIA [36] The MIN-K%++ is based on the assumption that, after tokenizing the sentence, the tokens that have been used for the training of the target model will be assigned with a higher confidence during the response compared to the possibility prediction of other tokens. To implement this assumption, assuming the whole sentence is $x=(x_{1},x_{2},\cdots,x_{t},\cdots,x_{N})$, it compared the possibility of the target token $x_{t}$ to the probability distribution of the whole vocabulary for the next token by following formula:
$MIN-K\%++(x_{1},x_{2},\cdots,x_{t})=\frac{\log p(x_{t}|x_{1\sim(t-1)})-\mu}{\sigma}$
The $\mu$ and $\sigma$ stand for the expectation and standard deviation of the probability distribution of the next token $x_{(t+1)}$’s log possibility. A higher score indicates that compared to the next token, the target model has higher confidence when calculating the current token, which indicates the current token may be presented in the training dataset and the next may not. After calculating the MIN-K%++ score of each token, similar to MIN-KPROB MIA [30], the system calculates the average of the kof the tokens that have the minimum score as the final score for the sentence. Depending on the pre-designed threshold, the identity could be predicted.
While the MIN-K%++ MIA uses the probability distribution of the next token to calibrate its score, it faces several limitations. Firstly, higher final scores indicate the consistency of the confidence of each token, which could indicate most of the tokens from the training dataset, so the sentence is used in training. However, if the sentence has most tokens not from the training dataset, the final score of that sentence could also be high.
Furthermore, a sentence containing most words presented in the training dataset doesn’t necessarily mean the sentence is in the training dataset.
3.2 Gray-Box & Reference Model Based
Likelihood Ratio MIA [28, 37] The strategy that supports likelihood ratio MIA is to use the likelihood ratio test to distinguish the scenario in which the target data appeared in the training dataset and the scenario in which the target data is independent. The likelihood of the former scenario can be calculated by querying the target model, which parameterized by $\theta$. For the latter scenario, a reference model trained on random data, which parameterized by ${\theta}_{r}$, is used to analyze the likelihood. The likelihood ratio test is implemented by the following equation, where s is the target data.
The prediction is based on comparing $\mathit{\Lambda}\left(s\right)$ to the threshold, which was generated by calculating the likelihood ratio score on general data with around 10% false positive rate tolerance.
Data Inference MIA [33] Unlike most MIA that use a single metric to judge the membership identity of the data, Data Inference MIA uses a collection of MIA methods for prediction. To train the attack model, the adversary has to collect a set of suspected data $D_{sus}$ and a set of data that is clearly not used during the training $D_{out}$. It is worth noticing that since $D_{out}$ can be easily generated, it didn’t require the collection of the member data. After collecting the data, the $D_{sus}$ and $D_{out}$ is further separate to $T_{sus}$, $V_{sus}$, $T_{out}$, and $V_{out}$. Using the MIA model including Min-kProb [30], Perplexity Based MIA [31], and several reference-based MIA, a feature map is drawn for each data. Using the data from $T_{sus}$ and $T_{out}$, an attack model is trained to distinguish the data from $D_{sus}$ and $D_{out}$. Then, the attacker uses $V_{sus}$ and $V_{out}$ to certify that this attack model is sufficient and the gap between $D_{sus}$ and $D_{out}$ do exist, which indicates the membership identity of $D_{sus}$.
Data Inference MIA eliminated the necessity of the pre-designed threshold by analyzing the existence of difference between the suspecter and outlier. However, depending on the data structure of the outlier collected, the system might be confused by the irrelevant differences and therefore have a high false-positive rate.
3.3 Gray-Box & Target model based
Semantic MIA [39] Following the sensitivity-based MIA, sematic MIA further controls the content of the perturbation added to input data to improve the attack performance. The attack is based on the similar assumption that the target model has different sensitivity for member and non-member data. However, to make sure the difference of the perturbation added each time does not disturb the attack, besides only considering the difference in prediction results between the original data and noisy data, the semantic distance brought by the perturbation, denote by $\phi\left(x\right)-\phi\left(x^{\prime}\right)$, is also analyzed. A neural network is used as the attack model $f_{attack\_NN}\left(x\right)$ and trained with supervised learning, which requires gray box access to get a portion of the training dataset.
Compared to sensitivity-based MIA, semantic MIA has dynamic adjustment over the different amounts of perturbation added. This approach approves the accuracy and robustness of the original method. The introduction of a neural network for the attack model enhances the capability of the attack model but brings the requirement of less restricted access.
Confidence-Based MIA [40] This Confidence-Based MIA utilizes the principle that target models usually have higher confidence when predicting data that it has seen. In LLM, this concept can be implemented by checking the confidence of the predicted current token given the preview tokens. Firstly, the MIA system attracts and normalizes the probability of each token from the input. Based on the calculated result, a feature map is generated for the input. The feature map will be fed into a classification model, which is trained by the data drawn from the training dataset and the independent data, to predict its membership information.
This MIA approach is less complex and requires less computation resources compared to some advanced algorithms. However, a huge difference in the confidence of member and non-member input is not guaranteed, which could lead to lower accuracy.
Data Inference-Based MIA [41] This MIA system is based on the data inference MIA [33]. Similarly, it deploys multiple MIA models to generate the feature map of each data collected, including perplexity-based MIA [31] and MIN-K%++ [36]. However, instead of training the attack model in the suspected dataset and outlier dataset, this system directly trains the attack model on pre-known member and non-member datasets, which was possible under the gray-box scenario.
Compared to the original attack method, the data inference-based MIA approach reduces the possibility that attack models are disturbed by irrelevant differences between the suspect dataset and outlier dataset and, therefore, have better performance. The drawback is requesting a higher level of access. The idea of embedding a variety of MIA methods makes these kinds of approaches more convincible compared to the others but may provide less contribution in the research dimension.
3.4 White-Box & Target Model Based
Compared to the origin sensitivity-based MIA and variance semantic MIA, noise neighbor achieves an even higher consistency in the perturbation added to the target data. The performance of the noise neighbors MIA model is comparable with the shadow model-based model but requires much fewer computation resources. However, the requirement of accessing the embedded layer may significantly limit its application.
Layer Activation-Based MIA The layer activation-based MIA systems predict membership information by examining the activation’s output of each neural. The LUMIA [43] implement this concept by adding Multi-Layer Perceptron (MLP), called Linear Probes (LP), to the output of each layer. LP generates its prediction in each layer, and LUMIA selects the highest score as final result.
The concept and implementation of the LUMIA are relatively simple and suitable for both LLMs and multimodal models. However, whether Multi-Layer Perceptron on the output of activations is sufficient enough to predict the membership information still needs further investigation. Only selecting the highest score may also make the system more vulnerable to bias or less accuracy.
3.5 Visualization of Taxonomy
To provide readers with a clear view and facilitate rapid navigation of the MIA algorithms being analyzed, Figure 1 presents a visualized categorization based on both attack scenarios and methodology.
[…]
my tree style/.style= for tree=grow=east, parent anchor=east, child anchor=west, where level=0my node=black,text width=7.5em,rotate=90,align=center,anchor=center, where level=1my node=black,text width=6em, where level=2my node=black,text width=6em, where level=3my node=black,text width=15em, l sep=1.5em, forked edge, fork sep=1em, edge=draw=black!50, thick, if n children=3for children= if n=2calign with current , tier/.option=level, {forest} my tree style [MIA against LLM [Methodology [Reference Model[Likelihood Ratio MIA[28, 37]
SPV-MIA[38]
Data Inference MIA[33]]] [Target Model[MIN-K%PROB[30]
Perplexity-Based MIA[31, 32]
Sensitivity-Based MIA[33, 34, 35]
MIN-K%++ MIA[36]
Semantic MIA[39]
Confidence-Based MIA[40]
Data Inference-Based MIA[41]
Layer Activation-Based MIA[43]
Noise Neighbor MIA[42]]] ] [Scenario [White-Box[Layer Activation-Based MIA[43]
Noise Neighbor MIA[42]]] [Gray-Box[Likelihood Ratio MIA[28, 37]
SPV-MIA[38]
Data Inference MIA[33]
Semantic MIA[39]
Confidence-Based MIA[40]
Data Inference-Based MIA[41]]] [Black-Box[MIN-K%PROB[30]
Perplexity-Based MIA[31, 32]
Sensitivity-Based MIA[33, 34, 35]
MIN-K%++ MIA[36]]] ] ]
Figure 1: Visualized Taxonomies for LLM focus MIA
4 Membership Inference Attack against Large Multimodal Model
4.1 Vision-language Model (VLM)
Feature-Based MIA The M4I MIA introduced by P. Hu et al. [44] is a feature-based MIA that explores the relation between the features of the images and texts pair, which is the input and output, for the prediction. Firstly, the attacker develops a neural model, including an image decoder for the feature extraction of the images, a text decoder for the analysis of the text input, and several fully connected layers for the fusion of encoded data, analysis of the relation between the features, and final classification. After that, under the gray box setting, the adversary trains the model with the member and non-member image-text pair. Finally, the attacker provides the attack model with the target input and the target model’s output to infer the identity.
For the limitation of this attack, the assumption that unique distinguishable feature exists between the image and text pair of the training dataset is not guaranteed.
Rényi entropy-based MIA [45] The MaxRényi MIA system utilizes the Rényi entropy to detect membership information. Similar to the other research, it assumes the target model has less perplexity when inferring data from the training dataset, which can be measured by less Rényi entropy. For the image MIA, the adversary first queries the target model with an image and promotion, with the description of the image returned. After that, the attacker queries the target with the image, promotion, and original response. The output logits from the target model are separate to image, instruction, and description logits. The final prediction can be made by calculating the Rényi entropy over the top Klogits. The pipeline introduced utilizes the instruction and description token instead of the image token, which makes it suitable for the attack in VLM since VLM has no image token. For the text MIA, since VLM do have text token, it can be achieved by computing the Rényi entropy over the top Kpositions.
Cosine Similarity-Based MIA [46] This attack algorithm is based on the assumption that the target model enhances the cosine similarity during the training. Therefore, the adversary can predict the identity based on the cosine similarity between the target input $x$ and its output $f_{target}\left(x\right)$, given by:
$CS\left(x,f_{target}\left(x\right)\right)>\tau.$
$CS\left(x,f_{target}\left(x\right)\right)-CS\left(T_{k}\left(x\right),f_{target}\left(x\right)\right)>\tau$
Unlike the original attack and AEA, which use cosine similarity directly for the prediction, Weakly Supervised Attack (WSA) uses it to collect meta-member datasets. The attacker first queries the target model with a dataset that surely is not used in training, $D_{out}$. The cosine similarity of the data and its answer from the $D_{out}$ follows the Gaussian distribution, allowing the attacker to decide a threshold at the higher end ($\mu+3\sigma$ for example) of the distribution. Such threshold is then applied to the general internet databases that might be used for the training to form a suspect member database $D_{sus}$. Finally, a binary classifier is trained to distinguish the feature of the image-text pair from $D_{out}$ and $D_{sus}$ as the attack model.
For the performance of the three approaches, AEA is marginally better than the base method, while WSA is generally better than AEA. However, WSA requires more computation resources and the knowledge of general databases that might be used for training. Furthermore, the assumption of maximizing cosine similarity may not always be valid depending on the specific algorithm of the target model.
Layer Activation-Based MIA The LUMIA [43] can also attack VLM with similar method mentioned in section 3.4.
Sensitivity-Based MIA The sensitivity-based MIA introduced by J. Ren et al. [34] in VLM is very similar to the one in LLM, which is based on the principle that member data is more sensitive and drops more when perturbation is added. Specifically, it paraphrases the text response as the perturbation and measures the average log-likelihood to justify the fluctuation. To design the threshold, it uses a part of the training dataset under the gray-box access and a reference dataset that was surely not used during the training to calculate the trend of the change of members and non-members at sample size.
Temperature-Based MIA The temperature-based MIA against VLM introduced by Y. Hu et al. [47] is based on the perplexity-based MIA that utilizes the accuracy of the target model for prediction. In their research, they further use temperature to magnify the difference in accuracy between member and non-member data. The research adapts this attack to multiple scenarios, including the situation only target data available, which compares the difference in the accuracy between the high temperature and low temperature with a pre-designed threshold, and the gray-box-like scenario by training the attack model based on the patterns of shadow-model’s members and non-members’ accuracy in different temperatures. Their research also contains a gray-box MIA Reference Inference MIA by comparing the target data with the reference set using the statistical hypothesis test (z-test), and a black-box Consistency-Based MIA that continues to query the target model and checks the consistency of the result with members tending to have more consistent outputs.
The research has analyzed multiple situations with different levels of understanding about the database, which increases the capability of the attack system. However, access to the temperature setting of the target model may already require white-box access since it is normally located in the SoftMax layer. The requirement may let the system less practical.
4.2 Multimodal Emotion Recognition (MER) Model
Representations-Based MIA [48] his white-box-based MIA model utilizes the representations, which are the encoded data of the sub-unimodal that analyzes a specific modality, of the input in each modality for the attack. A dense neural network binary classifier is used as the attack model. The adversary first collects a portion of the training dataset and a certain number of non-members, forming the attack dataset. The attacker fed the attack dataset to the target model to extract the representations of each data. These representations with the labels of member and non-member are used for the training of the attack model.
The attack algorithm is relatively easy to understand and requires less computational resources. However, without specification, a single neural network classifier might have limitations in fully finding the difference between member and non-member.
[…]
my tree style/.style= for tree=grow=east, parent anchor=east, child anchor=west, where level=0my node=black,text width=14em,rotate=90,align=center,anchor=center, where level=1my node=black,text width=6em, where level=2my node=black,text width=6em, where level=3my node=black,text width=15em, l sep=1.5em, forked edge, fork sep=1em, edge=draw=black!50, thick, if n children=3for children= if n=2calign with current , tier/.option=level, {forest} my tree style [MIA against Multimodal Model [Methodology [Reference Model[Temperature-Based MIA[47]]] [Target Model[Feature-Based MIA[44]
Rényi entropy-based MIA [45]Cosine Similarity-Based MIA[46]
Layer Activation-Based MIA[43]
Sensitivity-Based MIA[34]
Temperature-Based MIA[47]
Reference Inference MIA[47]
Consistency-Based MIA[47]
Representations-Based MIA[48]]] ] [Scenario [White-Box[Layer Activation-Based MIA[43]
Temperature-Based MIA[47]
Representations-Based MIA[48]]] [Gray-Box[Feature-Based MIA[44]
Cosine Similarity-Based MIA[46]
Sensitivity-Based MIA[34]
Cosine Similarity-Based MIA[46]
Consistency-Based MIA[47]]] ] [ Modality [MER[Representations-Based MIA[48]]] [VLM[Feature-Based MIA[44]
Rényi entropy-based MIA [45]
Cosine Similarity-Based MIA[46]
Layer Activation-Based MIA[43]
Sensitivity-Based MIA[34]
Temperature-Based MIA[47]
Reference Inference MIA[47]
Consistency-Based MIA[47]]] ] ]
Figure 2: Visualized Taxonomies for Large Multimodal Model focus MIA
4.3 Visualization of Taxonomy
5 Fine-tuning MIA in Advanced Machine Learning
The escalating scale and complexity of advanced machine learning models have propelled the adoption of fine-tuning techniques [41, 49], enabling the efficient deployment of these models across diverse domains and tasks. Unlike traditional methods that involve training models from scratch, fine-tuning leverages knowledge acquired during the pre-training phase, thereby achieving comparable performance with substantially reduced computational resources [50].
In the typical development pipeline, pre-training is normally conducted by model developers, while end-users or specific entities perform fine-tuning to tailor models to their unique applications. This pipeline lets fine-tuning often involves utilizing highly private or user-specific data. MIA aims to expose the data used during the training process. Its adoption in the fine-tuning process introduces an even heightened privacy concern.
In this section, we first refine the attack scenarios to account for the more restricted information accessibility during the fine-tuning stage. Following this, we review existing research efforts aimed at adapting MIAs to retrieve the fine-tuning datasets of LLMs and multimodal models.
5.1 Attack scenario
In the context of fine-tuning, our revised attack scenario definitions largely retain the conventional distinctions of black-box and white-box access, as outlined below:
Black-Box Access The adversary can only query the fine-tuned model, obtaining its inference results. Additionally, the attacker possesses general knowledge about the target model but lacks access to its internal parameters or training data.
White-Box Access The adversary has full access to both the pre-trained and fine-tuned models at the developer level. This includes complete knowledge of the model architecture, parameters, and both the pre-training and fine-tuning datasets.
The definition and characterization of gray-box access require refinement, as fine-tuning in real-world scenarios typically exhibits the following characteristics:

### v2

@@ Abstract
The adoption of the Large Language Model (LLM) has accelerated dramatically since ChatGPT from OpenAI went online in November 2022. Recent advances in Large Multimodal Models (LMMs), which process diverse data types and enable interaction through various channels, have expanded beyond the text-to-text limitations of early LLMs, attracting significant and concurrent attention from both researchers and industry. While LLMs and LMMs are starting to spread widely, concerns about their privacy risks are increasing as well. Membership Inference Attacks (MIAs) are techniques used to determine whether a particular data point was part of a model’s training set, which is a key metric for assessing the privacy vulnerabilities of machine learning models. Hu et al. show that various machine learning algorithms are vulnerable to MIA. Despite extensive studies on MIAs in classic models, there remains a lack of systematic surveys addressing their effectiveness and limitations in advanced large-scale models like LLMs and LMMs. In this paper, we systematically reviewed recent studies of MIA against LLMs and LMMs. We analyzed and categorized each attack based on its methodology, scenario, and targeted model, and we discussed the limitations of existing research. In addition to examining attacks on pre-training and fine-tuning stages, we also explore MIAs that target other development pipelines, including Retrieval-Augmented Generation (RAG) and the model alignment process. Based on the survey, we provide suggestions for future studies to improve the robustness of MIA in large-scale AI models.
@@ 1 Introduction
Pre-training and fine-tuning are two widely adopted training pipelines. A pre-trained model refers to a model that was trained by the original developer based on the pre-train dataset and starting from a random initial state. It usually shows an outstanding performance in general tasks, but may struggle to meet requirements under specified tasks. Fine-tuning is a technology that adapts a general pre-trained model to specific tasks. Unlike traditional methods that involve training models from scratch, fine-tuning leverages knowledge acquired during the pre-training phase, thereby achieving comparable performance with substantially reduced computational resources [11]. The escalating scale and complexity of advanced machine learning models have propelled the adoption of fine-tuning techniques [12, 13], enabling the efficient deployment of these models across diverse domains and tasks.
Several other pipelines may also being adapted to the machine learning to enhance its performance or meet specific requirements, including model alignment and Retrieval-Augmented Generation (RAG). Model alignment enhances models’ ability to generate responses that are more preferred by the user [14], which can be achieved by utilizing the human preference samples and techniques like Reinforcement Learning from Human Feedback (RLHF) [15]. To enhance models’ ability to generate more reliable and accurate responses, RAG stimulates the approach of citation and referencing [16]. Before the queries are actually fed into the model during the infer stage, the RAG system first modifies the original query. It searches through a reliable database and attaches relevant documentation to the query, which allows the model to generate responses referring to these trustworthy resources.
@@
As LLMs and LMMs continue to evolve and gain widespread adoption, addressing their privacy concerns has become increasingly critical. In the LLM area, various research studies have already been conducted that focus on the general privacy risks of LLM [23, 24]. However, these studies provide limited depth and focus in their analysis of MIA. In the field of multimodal models, apart from the work of M.A. Rahman et al. [25], which examines privacy concerns from an educational perspective, and a general review by S.K. Tetarave et al. [26], there remains a lack of in-depth and systematic analysis. Regarding MIA, the study conducted by H. Hu et al. [27] has summarized MIA against various basic machine learning models. While their survey provides a valuable reference for researchers in the area, the study is up to 2022, and there is a lack of study about MIA on more advanced machine learning models. A similar study was published in 2024 by J. Niu et al. [28], which still only summarizes the research attack up to Jan. 2024 and lacks analysis and focus on advanced models. Furthermore, there is still a lack of comprehensive review about MIA across different machine learning model development pipelines.Table 1 summarizes the limitations of existing surveys and highlights the contributions of our research.
@@
2. An analysis of MIA across different stages of the machine learning pipeline With the advancement of membership inference attack (MIA) techniques, their application has extended beyond the pre-training stage to encompass multiple phases of the machine learning development pipeline, like fine-tuning and alignment. To the best of our knowledge, this paper presents the first comprehensive survey that systematically analyzes MIA across these different stages.
3. Suggestions for future research. Based on the challenges and research gap identified, we provide recommendations to guide future studies in improving the robustness, applicability, and real-world relevance of MIA techniques..
@@
Research Large-Scale Model MIA Attack Across Survey Up
LLM LMM In-Depth Pipeline To
@@ 2 Adversarial Models in Membership Inference Attack
2.1 Targeted pipline
Pre-training MIA Pre-training is the first step in the training pipeline, and therefore is the original and most targeted part in MIAs. The pre-training MIA aims to expose the data from the pre-train dataset, which risks the privacy of the original model publisher.
Fine-tuning MIA In the typical development pipeline, model developers usually conduct pre-training, while end-users or specific entities perform fine-tuning to tailor models to their unique applications. This pipeline lets the fine-tuning process often involve utilizing highly private or user-specific data. As MIA aims to expose the data used during the training, its adoption in the fine-tuning process introduces an even heightened privacy concern.
Model alignment MIA Users’ Preference data is crucial for the model alignment process, which the user’s intention was learned based on. In model alignment MIA, the adversary aims to expose these sensitive data, which could lead to privacy leakage.
RAG MIA The RAG system utilizes a trustworthy database to provide the reference for the machine learning model during the infer stage. Such a database could be a critical and valuable asset for the developer and might also contain sensitive documentation. The RAD MIA aims to determine if a given target data is from this database.
@@ 2.2.1 General attack scenario
Table 2: Summarize of general attack scenario
@@
2.2.2 Fine-tuning MIA refinement
For the Fine-tuning MIA, since fine-tuning in real-world scenarios typically exhibits the following characteristics, the definition and characterization of gray-box access require refinement: